A number of inhibitors for studyase have been found, the following data from enzyme assay analysis has been determined:

S (mM)

Inhibitor A

V(μmole/mL/sec)

1mL of 1mM used

Inhibitor B

V(μmole/mL/sec)

1mL of 3mM used

Inhibitor C

V(μmole/mL/sec)

1mL of 0.05mM used

1

4.3

5.5

5

2

8

9

8.69

4

14

13

13.7

8

21

16

19.6

12

26

18

22.2

                                                     

Plot a Lineweaver-Burk plot of 1/V versus 1/S (on 1 graph) of each of the inhibitors.

Determine Vmax and Km for each inhibitor

Identify what type of inhibitor A,B and C are.

Calculate Kcat for each, do any of the inhibitors work better than the enzyme?

Would you expect any of the inhibitors to have a similar structure to studyase? Why?
Answer #1

0.25 y 0.2118x+0.02 R2 0.9998 0.2 y 0.1695x+0.0304 | y-01372x + 00438 R20.9994 0.15 1/C Lineal (1/A) ....Lineal (1/c) Lineal

in order to determinate the kind of inhibitor it is need the values without inhibitor because we need to compare

the cut in y-axis is 1/Vmax, and the cut in x-axis is -1/km

for...

A) y = 0.2118x+0.02

x=0-> y = 0.02 : Vmax = 1/0.02 = 50

0=0.2118x+0.02 :x = -0.094429 : Km = 10.59

B) Vmax = 32.89, km= -5.57

C) Vmax = 22.83, km = 3.13

Please post the values without inhibition in order to solve the other questions